OpenStudy (anonymous):

How to find the right and left limit of lim x-->3 (sqrt(x)-sqrt(3))/(x-3) +3x^2 ?

OpenStudy (anonymous):

\[\lim_{x \rightarrow 3} (\sqrt{x}-\sqrt{3})\div( x-3) + 3x^{2}\]

OpenStudy (mathmale):

Please clarify: are you dividing by [ (x-3) + 3x^2 ] or only by (x-3). If the former, then please make that clear by using another set of parentheses to enclose the denominator. This math problem is from what course?

OpenStudy (anonymous):

\[\lim_{x\to3}\left(\frac{\sqrt x-\sqrt3}{x-3}+3x^2\right) ~~?\]

OpenStudy (anonymous):

oops sorry! yes, what sithsandgiggles posted is how the equation should be set up.

OpenStudy (anonymous):

and this is from calc 1.

OpenStudy (mathmale):

You'll get faster (and perhaps better) responses from potential helpers if you'd please be certain that the expression you type in is in no way ambiguous.

OpenStudy (mathmale):

As x approaches 3, 3x^2 approaches 27. That's the easy part. As for the first term within parentheses: Try factoring x-3. Hint: Note that Sqrt(x)-Sqrt(3) is one of the factors of x-3.

OpenStudy (anonymous):

So could I factor it as \[(\sqrt{x}-\sqrt{3})(\sqrt{x}-\sqrt{3})\] to get x-3?

OpenStudy (anonymous):

oops, i meant to put a plus instead of a minus for one of the factors.